OpenStudy (anonymous):

Help Solve by elimination. -4x +5y=14 3x +y=20

OpenStudy (anonymous):

[-4x+5y=14]* 3 [3x+y=20] * -4 ________________ -12x+15y=42 -12x-4x=-80 _______________- 19y=122 y=122/19

OpenStudy (anonymous):

Hey thanks - it make more sense now.

OpenStudy (anonymous):

No problem. 122/19 seems like a weird answer though so I'd double check. You can also plug in your y value into either equation to solve for x

OpenStudy (anonymous):

y=11.09 repeating, x=10.36 repeating is what I come up with